# On the Parallel Complexity of Identifying Groups and Quasigroups via Decompositions

### v1

@@ 4.2 Conjugacy Class Graph of a Group
In this section, we will leverage the conjugacy class of a group (to be defined shortly) to compute a fully-refined direct product decomposition of a non-Abelian group. We first recall key preliminaries from [KN09, Section 6]. Let $G$ be a group. For $g\in G$, let $\mathcal{C}_{g}$ be the conjugacy class of $g$. We say that two conjugacy classes $\mathcal{C}_{g},\mathcal{C}_{h}$ of $G$ commute if for all $\alpha\in\mathcal{C}_{g}$ and all $\beta\in\mathcal{C}_{h}$, $\alpha$ and $\beta$ commute.
@@
Let $G=G_{1}\times\cdots\times G_{k}$. If a conjugacy class $\mathcal{C}_{g}$ is irreducible, then there exists a unique $i\in[k]$ such that $\pi_{i}(k)\not\in Z(G_{i})$, where $\pi_{i}:G\to G_{i}$ is the canonical projection map.
@@
Let $G$ be a group. We can compute $\Gamma_{G}$ in L. Furthermore, we can compute the connected components of $\Gamma_{G}$ in L.
@@
First, we observe that we can compute the conjugacy classes of $G$ in $\mathsf{AC^{0}}$. We now claim that for a given conjugacy class $\mathcal{C}_{x}$, we can decide if $\mathcal{C}_{x}$ is irreducible in L. In $\mathsf{AC^{0}}$, we can test whether $x\in Z(G)$. Furthermore, in $\mathsf{AC^{0}}$, we may examine all pairs $a,b\in G$ and test whether $\mathcal{C}_{x}=\mathcal{C}_{a}\cdot\mathcal{C}_{b}$. Now in L, we may test whether $|\mathcal{C}_{x}|=|\mathcal{C}_{a}|\cdot|\mathcal{C}_{b}|$. Thus, in L, we may identify the irreducible conjugacy classes of $G$.
We now turn to writing down $\Gamma_{G}$. The irreducible conjugacy classes of $G$ are precisely the vertices of $\Gamma_{G}$. Now let $\mathcal{C}_{a},\mathcal{C}_{b}$ be two such irreducible conjugacy classes. The vertices corresponding to $\mathcal{C}_{a},\mathcal{C}_{b}$ are adjacent if and only if for every $g\in\mathcal{C}_{a}$ and every $h\in\mathcal{C}_{b}$, $gh=hg$. Thus, we can decide in $\mathsf{AC^{0}}$ whether $\mathcal{C}_{a},\mathcal{C}_{b}$ are adjacent in $\Gamma_{G}$. Once we have $\Gamma_{G}$, we may compute the connected components in L [Rei08]. The result now follows. ∎
@@ 4.3 Algorithm
We now parallelize the main direct product decomposition algorithm of Kayal and Nezhmetdinov [KN09, Algorithm 1]. Let $G$ be our input group. If $G$ is Abelian, then we can decompose $G$ in $\textsf{AC}^{1}$ using Proposition 4.5. So suppose $G$ is non-Abelian. We begin by computing the conjugacy class graph $\Gamma_{G}$, and the connected components $\Lambda_{1},\ldots,\Lambda_{t}$ of $\Gamma_{G}$. By Lemma 4.11, this step is L-computable. Now if $G$ admits a non-trivial direct product decomposition, then there exists some $S_{1}\subseteq[t]$ inducing an indecomposable direct factor of $G$ [KN09]. By Proposition 4.12, we have that $t\leq\log n$. Thus, we may try all such $S_{1}$ in parallel.
@@
We may write down the elements of $\text{Elts}(\Lambda_{j})$ ($j\not\in S_{1}$) in $\mathsf{AC^{0}}$. We may then use a membership test to, in L, write down:
@@
Given $X$, we may write down $Z_{1}=C_{G}(X)$ in $\mathsf{AC^{0}}$. Thus, we may write down $Z_{1}$ in L. Using Proposition 4.7, we can in $\textsf{AC}^{2}$, determine $H_{1},Y\trianglelefteq G$ such that $Z_{1}=H_{1}\times Y$, where $H_{1}$ has no Abelian direct factors and $Y$ is Abelian. We now use Proposition 4.6 to decide if there exists $Y_{1}\trianglelefteq G$ such that $G=H_{1}\times Y_{1}$. If no such $Y_{1}$ exists, then $H_{1}$ is not an indecomposable direct factor of $G$. Otherwise, we recursively compute a direct product decomposition of $Y_{1}$.
@@ 5 Canonizing Direct Products in Parallel
In order to obtain a canonical labeling, we individualize each non-identity element from $g_{1},\ldots,g_{d}$ such that for all distinct $i,j\in[d]$, $g_{i}$ and $g_{j}$ receive different colors. We then run $(d,O(1))$-WL, which we again note is L-computable. Each color class will now have size $1$. We label an element $g\in G$ according to the color class of the $d$-tuple $(g,\ldots,g)$.
@@
Now we run $(d,0)$-WL Version II on $G$. For each $i\in[k]$, we may in $\mathsf{AC^{0}}$, identify the colored $d$-tuples of $G_{i}^{d}$. As only the initial coloring of WL has occurred, observe that the partition on $G_{i}^{d}$ obtained by these colored $d$-tuples agrees exactly with that of applying $(d,0)$-WL Version II to $G_{i}$. Thus, we may apply Lemma 5.1 to obtain a canonical labeling $\lambda_{i}$ of each $G_{i}$ and a canonical generating sequence $(g_{i,1},\ldots,g_{i,d})$ in L. In particular, Lemma 5.1 guarantees that $\chi_{d,0}((g_{i,1},\ldots,g_{i,d}))$ is the minimum color class amongst all such generating sequences in $G_{i}^{d}$.
We now sort $G_{1},\ldots,G_{k}$ such that $G_{i}$ appears before $G_{j}$ whenever $\chi_{d,0}((g_{i,1},\ldots,g_{i,d}))<\chi_{d,0}((g_{j,1},\ldots,g_{j,d}))$. By re-indexing $G_{1},\ldots,G_{k}$, we may assume (without loss of generality) that the direct factors appear in order $(G_{1},\ldots,G_{k})$. This step is L-computable. In order to obtain a canonical labeling, we now individualize the non-identity elements of our canonical tuples (without loss of generality, we may assume that all such elements are not the identity):
@@
such that for all distinct pairs $(i,j),(i^{\prime},j^{\prime})\in[k]\times[d]$, $g_{i,j}$ and $g_{i^{\prime},j^{\prime}}$ are individualized to receive a different color. We then run $(O(1),O(1))$-WL Version II on the original uncolored group $G$, which is L-computable. Observe that for each $i\in[k]$, the resulting coloring induces the labeling $\lambda_{i}$ on $G_{i}$, and hence subsequently assigns each element in $G$ a unique color. Precisely, we assign to an element $g\in G$ the label corresponding to the color class $\chi_{O(1),O(1)}((g,\ldots,g))$.
@@ 6.2 Pointwise Transporter on Small Domains in Parallel
In the permutation group model, a group $G$ is specified succinctly by a sequence of permutations from $\text{Sym}(m)$. The computational complexity for this model will be measured in terms of $m$. We will recall the following standard suite of NC algorithms in the setting of permutation groups.
@@
The Pointwise Transporter problem takes as input $G\leq\text{Sym}(m)$ given by a sequence of generators, and two sequences $(x_{1},\ldots,x_{k})$, $(y_{1},\ldots,y_{k})$, and tests whether there exists some $g\in G$ such that for all $i\in[k]$, $x_{i}^{g}=y_{i}$. We will establish the following:
@@ 6.3 Isomorphism Testing of Central Quasigroups
Let (Q,)(Q,*) be a quasigroup, and fix $e\in Q$ arbitrarily. We can write down the set $D=\{\lambda_{x}\lambda_{e}^{-1}:x\in Q\}\subseteq\mathrm{Dis}(Q,*)$ and check if $\mathrm{Dis}(Q,*)$ is isotopic to a group; and if so, then we can also write down the multiplication table for $\mathrm{Dis}(Q,*)$ in $\mathsf{AC^{0}}$.
@@
Let (Q,)(Q,*) be a quasigroup. We can decide in $\textsf{AC}^{0}$ whether $G$ is central and construct its representation $\mathcal{Q}(Q,+,\phi,\psi,c)$.
@@
By Proposition 6.9, we can decide in $\textsf{AC}^{0}$ whether both $G_{1},G_{2}$ are central; and if so, construct the Abelian groups $\mathrm{Dis}(G_{i},*_{i})$ and their representations $\mathcal{Q}_{i}(G_{i},+,\phi_{i},\psi_{i},c_{i})\cong G_{i}$. Now using Proposition 4.5, we may in $\textsf{AC}^{1}$ compute a basis $b_{i1},\ldots,b_{ik}$ for $\mathrm{Dis}(G_{i},*_{i})$, and sort the elements so that $|b_{ij}|\leq|b_{i,j+1}|$ for all $i,j$. Note that $k\leq\log|G_{i}|$. In particular, this allows us to decide if $\mathrm{Dis}(G_{1},*_{i})\cong\mathrm{Dis}(G_{2},*_{2})$ are isomorphic; and if so, construct an isomorphism $\tau:\mathrm{Dis}(G_{1},*_{1})\cong\mathrm{Dis}(G_{2},*_{2})$. Let $(G,+)$ be the Abelian group such that $(G,+)\cong\mathrm{Dis}(G_{i},*_{i})$.
@@ 7 Conclusion
Note that $\mathcal{D}$ includes the class of all Abelian groups. Grochow and Levet [GL23b] showed that the class of all Abelian groups has count-free WL dimension $\Theta(\log n)$.
@@
Grochow and Levet [GL23b] showed that $(O(1),O(\log n))$-WL Version II identifies $\mathcal{D}$. It is unclear how to approach this, even in the restricted case when each group admits a unique fully-refined direct product decomposition.
@@
In the setting of groups specified by their multiplication tables, can we compute fully-refined direct product decompositions in L?
@@
$\forall^{\log\log n}\textsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$, which is a proper subclass of L [CGLW24]. As central quasigroups are precisely the Abelian objects (in the sense of universal algebra) in the variety of quasigroups [Sze98], L is a natural target.

### v2

@@ 4.2 Conjugacy Class Graph of a Group
In this section, we will leverage the conjugacy class graph of a group (to be defined shortly) to compute a fully-refined direct product decomposition of a non-Abelian group. We first recall key preliminaries from [KN09, Section 6]. Let $G$ be a group. For $g\in G$, let $\mathcal{C}_{g}$ be the conjugacy class of $g$. We say that two conjugacy classes $\mathcal{C}_{g},\mathcal{C}_{h}$ of $G$ commute if for all $\alpha\in\mathcal{C}_{g}$ and all $\beta\in\mathcal{C}_{h}$, $\alpha$ and $\beta$ commute.
@@
Let $G=G_{1}\times\cdots\times G_{k}$. If a conjugacy class $\mathcal{C}_{g}$ is irreducible, then there exists a unique $i\in[k]$ such that $\pi_{i}(g)\not\in Z(G_{i})$, where $\pi_{i}:G\to G_{i}$ is the canonical projection map.
@@
Let $G$ be a group. We can compute $\Gamma_{G}$ in FL. Furthermore, we can compute the connected components of $\Gamma_{G}$ in FL.
@@
First, we observe that we can compute the conjugacy classes of $G$ in $\mathsf{AC^{0}}$. We now claim that for a given conjugacy class $\mathcal{C}_{x}$, we can decide if $\mathcal{C}_{x}$ is irreducible in L. In $\mathsf{AC^{0}}$, we can test whether $x\in Z(G)$. Furthermore, in $\mathsf{AC^{0}}$, we may examine all pairs $a,b\in G$ and test whether $\mathcal{C}_{x}=\mathcal{C}_{a}\cdot\mathcal{C}_{b}$. Now in L, we may test whether $|\mathcal{C}_{x}|=|\mathcal{C}_{a}|\cdot|\mathcal{C}_{b}|$. Thus, in FL, we may identify the irreducible conjugacy classes of $G$.
We now turn to writing down $\Gamma_{G}$. The irreducible conjugacy classes of $G$ are precisely the vertices of $\Gamma_{G}$. Now let $\mathcal{C}_{a},\mathcal{C}_{b}$ be two such irreducible conjugacy classes. The vertices corresponding to $\mathcal{C}_{a},\mathcal{C}_{b}$ are adjacent if and only if for every $g\in\mathcal{C}_{a}$ and every $h\in\mathcal{C}_{b}$, $gh=hg$. Thus, we can decide in $\mathsf{AC^{0}}$ whether $\mathcal{C}_{a},\mathcal{C}_{b}$ are adjacent in $\Gamma_{G}$. Once we have $\Gamma_{G}$, we may compute the connected components in FL [Rei08]. The result now follows. ∎
@@ 4.3 Algorithm
We now parallelize the main direct product decomposition algorithm of Kayal and Nezhmetdinov [KN09, Algorithm 1]. Let $G$ be our input group. If $G$ is Abelian, then we can decompose $G$ in $\textsf{AC}^{1}$ using Proposition 4.5. So suppose $G$ is non-Abelian. We begin by computing the conjugacy class graph $\Gamma_{G}$, and the connected components $\Lambda_{1},\ldots,\Lambda_{t}$ of $\Gamma_{G}$. By Lemma 4.11, this step is FL-computable. Now if $G$ admits a non-trivial direct product decomposition, then there exists some $S_{1}\subseteq[t]$ inducing an indecomposable direct factor of $G$ [KN09]. By Proposition 4.12, we have that $t\leq\log n$. Thus, we may try all such $S_{1}$ in parallel.
@@
We may write down the elements of $\text{Elts}(\Lambda_{j})$ ($j\not\in S_{1}$) in $\mathsf{AC^{0}}$. We may then use a membership test to, in FL, write down:
@@
Given $X$, we may write down $Z_{1}=C_{G}(X)$ in $\mathsf{AC^{0}}$. Thus, we may write down $Z_{1}$ in FL. Using Proposition 4.7, we can in $\textsf{AC}^{2}$, determine $H_{1},Y\trianglelefteq G$ such that $Z_{1}=H_{1}\times Y$, where $H_{1}$ has no Abelian direct factors and $Y$ is Abelian. We now use Proposition 4.6 to decide if there exists $Y_{1}\trianglelefteq G$ such that $G=H_{1}\times Y_{1}$. If no such $Y_{1}$ exists, then $H_{1}$ is not an indecomposable direct factor of $G$. Otherwise, we recursively compute a direct product decomposition of $Y_{1}$.
@@ 5 Canonizing Direct Products in Parallel
In order to obtain a canonical labeling, we individualize each non-identity element from $g_{1},\ldots,g_{d}$ such that for all distinct $i,j\in[d]$, $g_{i}$ and $g_{j}$ receive different colors. We then run $(d,O(1))$-WL, which we again note is FL-computable. Each color class will now have size $1$. We label an element $g\in G$ according to the color class of the $d$-tuple $(g,\ldots,g)$.
@@
Now we run $(d,0)$-WL Version II on $G$. For each $i\in[k]$, we may in $\mathsf{AC^{0}}$, identify the colored $d$-tuples of $G_{i}^{d}$. As only the initial coloring of WL has occurred, observe that the partition on $G_{i}^{d}$ obtained by these colored $d$-tuples agrees exactly with that of applying $(d,0)$-WL Version II to $G_{i}$. Thus, we may apply Lemma 5.1 to obtain a canonical labeling $\lambda_{i}$ of each $G_{i}$ and a canonical generating sequence $(g_{i,1},\ldots,g_{i,d})$ in FL. In particular, Lemma 5.1 guarantees that $\chi_{d,0}((g_{i,1},\ldots,g_{i,d}))$ is the minimum color class amongst all such generating sequences in $G_{i}^{d}$.
We now sort $G_{1},\ldots,G_{k}$ such that $G_{i}$ appears before $G_{j}$ whenever $\chi_{d,0}((g_{i,1},\ldots,g_{i,d}))<\chi_{d,0}((g_{j,1},\ldots,g_{j,d}))$. By re-indexing $G_{1},\ldots,G_{k}$, we may assume (without loss of generality) that the direct factors appear in order $(G_{1},\ldots,G_{k})$. This step is FL-computable. In order to obtain a canonical labeling, we now individualize the non-identity elements of our canonical tuples (without loss of generality, we may assume that all such elements are not the identity):
@@
such that for all distinct pairs $(i,j),(i^{\prime},j^{\prime})\in[k]\times[d]$, $g_{i,j}$ and $g_{i^{\prime},j^{\prime}}$ are individualized to receive a different color. We then run $(O(1),O(1))$-WL Version II on the original uncolored group $G$, which is FL-computable. Observe that for each $i\in[k]$, the resulting coloring induces the labeling $\lambda_{i}$ on $G_{i}$, and hence subsequently assigns each element in $G$ a unique color. Precisely, we assign to an element $g\in G$ the label corresponding to the color class $\chi_{O(1),O(1)}((g,\ldots,g))$.
@@ 6.2 Pointwise Transporter on Small Domains in Parallel
In the permutation group model, a group $G$ is specified succinctly by a generating sequence of permutations from $\text{Sym}(m)$. The computational complexity for this model will be measured in terms of $m$. We will recall the following standard suite of NC algorithms in the setting of permutation groups.
@@
The Pointwise Transporter problem takes as input $G\leq\text{Sym}(m)$ given by a sequence of generators, and two sequences $(x_{1},\ldots,x_{k})$, $(y_{1},\ldots,y_{k})$ of points from $[m]$, and tests whether there exists some $g\in G$ such that for all $i\in[k]$, $x_{i}^{g}=y_{i}$. We will establish the following:
@@ 6.3 Isomorphism Testing of Central Quasigroups
Let (Q,)(Q,*) be a quasigroup, and fix $e\in Q$ arbitrarily. We can write down the set $D=\{\lambda_{x}\lambda_{e}^{-1}:x\in Q\}\subseteq\mathrm{Dis}(Q,*)$ and check if (Q,)(Q,*) is isotopic to a group; and if so, then we can also write down the multiplication table for $\mathrm{Dis}(Q,*)$ in $\mathsf{AC^{0}}$.
@@
Let (Q,)(Q,*) be a quasigroup. We can decide in $\textsf{AC}^{0}$ whether (Q,)(Q,*) is central and construct its representation $\mathcal{Q}(Q,+,\phi,\psi,c)$.
@@
By Proposition 6.9, we can decide in $\textsf{AC}^{0}$ whether both $G_{1},G_{2}$ are central; and if so, construct the Abelian groups $(G_{i},+)\cong\mathrm{Dis}(G_{i},*_{i})$ and their representations $\mathcal{Q}_{i}(G_{i},+,\phi_{i},\psi_{i},c_{i})\cong G_{i}$. Now using Proposition 4.5, we may in $\textsf{AC}^{1}$ compute a basis $b_{i1},\ldots,b_{ik}$ for $\mathrm{Dis}(G_{i},*_{i})$, and sort the elements so that $|b_{ij}|\leq|b_{i,j+1}|$ for all $i,j$. Note that $k\leq\log|G_{i}|$. In particular, this allows us to decide if $\mathrm{Dis}(G_{1},*_{i})\cong\mathrm{Dis}(G_{2},*_{2})$ are isomorphic; and if so, construct an isomorphism $\tau:\mathrm{Dis}(G_{1},*_{1})\cong\mathrm{Dis}(G_{2},*_{2})$. Let $(G,+)$ be the Abelian group such that $(G,+)\cong\mathrm{Dis}(G_{i},*_{i})$.
@@ 7 Conclusion
Note that $\mathcal{D}$ includes the class of all Abelian groups. Grochow and Levet [GL26] showed that the class of all Abelian groups has count-free WL dimension $\Theta(\log n)$.
@@
Grochow and Levet [GL26] showed that $(O(1),O(\log n))$-WL Version II identifies $\mathcal{D}$. It is unclear how to approach this, even in the restricted case when each group admits a unique fully-refined direct product decomposition.
@@
In the setting of groups specified by their multiplication tables, can we compute fully-refined direct product decompositions in FL?
@@
$\forall^{\log\log n}\textsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$, which is a proper subclass of L [CGLW25]. As central quasigroups are precisely the Abelian objects (in the sense of universal algebra) in the variety of quasigroups [Sze98], L is a natural target.